OpenStudy (anonymous):

What is y-3=3(x+1) in standard form of a perpendicular line that passes through (5,1)?

OpenStudy (anonymous):

if m is the slope of a line then slope of a line perpendicular to the line is -1/m

OpenStudy (anonymous):

y-3=3(x+1) y= 3x+3+3 y= 3x+6 comparing it with y=mx+c we find slope m= 3 therefore slope of line perpendicular to given line i.e. \[m _{1}=-\frac{1}{3}\] Now eq of required line in slope point form is given as: \[y-y_{1}=m _{1}(x-x_{1})\] where x1=5 anf y1= 1 thus \[y-1=-\frac{1}{3}(x-5) \rightarrow 3y-3=-x+5\rightarrow x+3y =8 \] Hence x+3y =8 is the required of the line

OpenStudy (anonymous):

What is the X-intercept of the perpendicular line?